OpenStudy (anonymous):

Andrew can paint the neighbor's house 6 times as fast as Bailey. The year Andrew and Bailey worked together, it took them 1 days. How long would it take each to paint the house?

OpenStudy (anonymous):

This is exactly how the question is asked.. It makes no sense to me...

OpenStudy (anonymous):

A = andrew's painting rate B = bailey's painting rate H = house size H/(A+B) = 1 A = 6B

OpenStudy (anonymous):

Im still confused.

OpenStudy (anonymous):

Does what I've written so far make sense to you?

OpenStudy (anonymous):

Not really.I mean I understand the A and B and H.. But putting them together to h/(A+B)=1, I don't understand.

OpenStudy (anonymous):

(A+B) is how quickly they can paint together. H is how much they have to paint. So H/(A+B) gives the time it takes to paint the house. We know that time is 1. So H/(A+B) = 1.

OpenStudy (anonymous):